OpenStudy (anonymous):

lines v and w are parallel.which statement can you not deduce from this information

OpenStudy (anonymous):



OpenStudy (anonymous):

Choices?

OpenStudy (anonymous):

<5 is congruent ti 8 <1is congruent to <5 m<2+m<6=180 m<3+m<5=180

OpenStudy (anonymous):

m<2+m<6=180

OpenStudy (anonymous):

great help